# Shortcomings of LLMs for Low-Resource Translation: Retrieval and Understanding are Both the Problem

### v1

@@ Abstract
This work investigates the in-context learning abilities of pretrained large language models (LLMs) when instructed to translate text from a low-resource language into a high-resource language as part of an automated machine translation pipeline. We conduct a set of experiments translating Southern Quechua to Spanish and examine the informativity of various types of information retrieved from a constrained database of digitized pedagogical materials (dictionaries and grammar lessons) and parallel corpora. Using both automatic and human evaluation of model output, we conduct ablation studies that manipulate (1) context type (morpheme translations, grammar descriptions, and corpus examples), (2) retrieval methods (automated vs. manual), and (3) model type. Our results suggest that even relatively small LLMs are capable of utilizing prompt context for zero-shot low-resource translation when provided a minimally sufficient amount of relevant linguistic information. However, the variable effects of prompt type, retrieval method, model type, and language-specific factors highlight the limitations of using even the best LLMs as translation systems for the majority of the world’s 7,000+ languages and their speakers.
@@ 1 Introduction
The field has made great progress improving the quality of machine translation (MT) systems, but constraints on the amount and kinds of data available in the majority of the world’s 7,000+ languages have led to yet another disparity in access and support for speakers of these languages: low-resource MT continues to be a major challenge Hendy et al. (2023); Nicholas and Bhatia (2023); Robinson et al. (2023); Stap and Araabi (2023). While many of these languages lack the kinds of large, standardized corpora necessary for traditional methods, recent work shows it may be possible to leverage a smaller amount of existing resources, for example pedagogical materials used for language instruction, with Large Language Models (LLMs), albeit with varying results Elsner and Needle (2023); Tanzer et al. (2024); Zhang et al. (2024). These materials are often the result of community-driven or government-led initiatives to support language revitalization, reclamation, and mother-tongue education Schreiner et al. (2020); Liu et al. (2022); Riestenberg et al. (2024). Such discrepancies in the needs and priorities of academic, commercial, and community-led efforts to develop digital resources and language technologies is what Gessler (2022) terms the “NLP Gap”.
@@
Our results suggest that while, unsurprisingly, translation quality improves with model size, such improvements seem to primarily be the result of previous exposure to the low-resource language during model pretraining, rather an improved ability for the model to utilize prompt context, as evidenced by high scores in response to baseline (zero-shot) translation prompts. However, we also find evidence that in-context learning abilities may be inconsistent across different models of similar size. As found in previous studies, prompts containing morpheme and word-level translations reliably improve model outputs, but information from the grammar and corpus have a null or even negative effect on results. Human evaluation on a selection of outputs from two models – GPT-3.5 Turbo and GPT-4o – align with the quantitative measures we obtain using BLEURT Sellam et al. (2020) as an automatic metric. Quantitative results also show an effect of automated retrieval on translation quality that is most evident in prompts containing morpheme translations and for models with lower baseline scores. Finally, we highlight a number of ethical concerns and limitations that arise from the proposed methods that are supported by our findings, and discuss the potential risks and challenges LLM-based methods for low-resource MT face moving forward.
@@ 2 LLMs for Machine Translation
Zhang et al. (2024) find that sentences from the corpus retrieved using BM25 embeddings Robertson et al. (2009) work better than random ones. Tanzer et al. (2024), however, report that retrieval with longest common substring (LCS) matching outperforms embedding-based retrieval. Overall, the question of how to best retrieve relevant passages containing grammar material or sentences in a low-resource language is still open. This also complicates the interpretation of the mostly-negative results found for grammar passages. It is not clear whether these stem from poor retrieval, from the LLMs’ inability to process the retrieved content, or both. Moreover, although Reid et al. (2024) conducts human evaluation of the results for quality, to the best of our knowledge no study to date systematically investigates specific grammatical errors in the output.
@@ 3 Quechuan Languages
Quechua is a family of languages indigenous to the Andes in South America. This study focuses on varieties of Southern Quechua (S. Quechua, also known as urin quechua or quechua sureño) spoken in parts of Peru. [^1] Unless noted otherwise, we use Quechua in this study to refer Southern Quechua and related varieties, following the practices of native speakers with whom we have relationships. While previous studies investigated language/LLM pairs for which the baseline LLM lacked any pretrained knowledge, we find that newer LLMs can translate some S. Quechua sentences in a zero-shot setting. We expect this to be typical of many low-resource languages which, while often endangered, still may have some presence on the web.
@@ 4.1 Data
We conduct experiments on a collection of 50 pairs of S. Quechua - Spanish sentences sourced from one of the author’s personal notes. These were selected to highlight a range of specific grammatical phenomena at multiple levels of difficulty— they include simple clauses and tenses (e.g., qam allinta tusunki (tu bailas bien) ‘you dance well’) as well as more advanced constructions such as those involving past participles (e.g., awasqay waliqa sumaqmi (la falda que tejí es linda) ‘the skirt I knit is lovely’ and simultaneous events (e.g., qamqa takita uyarispa wasiykita pichachkanki (tú estás limpiando tu casa escuchando música) ‘you’re cleaning your house listening to music’. The first author, a foreign-language student of S. Quechua, received permission from her instructor to use notes from their lessons for the study. All sentence pairs were inspected by the instructor, a native bilingual speaker of both S. Quechua and Peruvian Spanish, to eliminate any errors and confirm the accuracy of all reference translations.
@@ 4.2 Prompt Construction
quechua: kay wasiqa turiypam
español:
Figure 1: Example baseline prompt. English: [TASK] Translate the following sentence from Quechua to Spanish. Respond only with the translation: Quechua: kay wasiqa tirypam; Spanish:
@@ 4.2.1 Morpheme Translations (morph)
We use a morphological parser Rios (2015) to segment each word of the source segment into canonical morphemes, each with gloss symbols and a Spanish translation. [^2] We set aside valid concerns regarding the theoretical status of the morpheme for this study and define a morph(eme) loosely as a recognizable form-meaning pair that recurs in a language. Some morphemes have multiple candidate meanings, all of which are retrieved. As an example, the word rantikuq is segmented as ranti-ku-q and glossed as “comprar.DB.VRoot-DB.VDeriv.+RflxInt-+Ag.NS.” While numerous orthographic standards have been developed and promoted across Quechuan-speaking communities in South America, considerable variation in orthographic conventions may be found even within a particular community or variety Rios and Castro Mamani (2014). We discuss the implications of this for our results in Section 4.2.5.
We supplement the output from the parser using a Quechua-Spanish bilingual dictionary Qheswa Simi Hamut’ana Kurak Suntur (2005). We retrieve any dictionary entry whose headword exactly matches a canonical morpheme in our segmentation. By default, we include all senses and any usage examples or contextual information in the dictionary entry as part of the prompt. We then concatenate the output of the parser with the retrieved dictionary entries and include this morph information as prompt context preceding the source sentence and baseline translation prompt.
@@ 4.2.2 Grammar Descriptions (grammar)
We also experiment with the inclusion of grammar lessons found in student-facing pedagogical materials, retrieving grammatical explanations relevant to each source sentence from a PDF document developed for students and teachers of S. Quechua. Pinto Tapia et al. (2005). The document is organized into short sections (1-3 sentences, plus paradigm tables or usage examples) that describe the particular grammatical concept associated with an affix in Quechua. For each source sentence, we retrieve sections associated with any affix listed in the document that is an exact match of a canonical morpheme and include this in prompts using contextual information from the grammar. This improves on the methods described in Tanzer et al. (2024), who use LCS-based retrieval over an entire textbook, and Elsner and Needle (2023); Zhang et al. (2024), whose grammatical description remains consistent across prompts regardless of the source text being translated.
@@ 4.4 Evaluation
We conduct both automatic and human evaluations to identify trends in model errors and outputs in the various experimental conditions. We use BLEURT as an automatic metric, and report mean BLEURT scores across items as the primary quantitative measure of translation quality for each of the conditions and models. We also use an adapted MQM schema to conduct qualitative human evaluation of the outputs of GPT-3.5 and GPT-4o for all prompts with automatic retrieval.
Each item selected for human evaluation is annotated by at least one of the authors by comparing the model’s output to the source text and reference translation. [^4] We discuss limitations on this process given the authors’ respective proficiency levels in S. Quechua and Spanish, as well as the steps we take to address them, in Section 8. We refer to the complete MQM typology to design our own four-dimensional framework of commonly attested errors in LLM-MT, each with a defined set of specific subtypes. Precise definitions and examples for all error categories and subtypes may be found in Appendix D.
Many of the categories in our schema are defined as in the core MQM framework. However, to capture some of the key behaviors reported in previous studies on LLM-MT and to evaluate the effects of prompt type on model outputs, we make the following adjustments. First, we utilize the Addition and Omission errors defined as Accuracy subtypes in the original MQM typology, but distinguish these from three additional subtypes: Substitution - Incorrect Subject, Substitution - Incorrect Tense/Aspect/Modality (TAM), and Substitution - Other. This is intended to capture LLM translations that differ from the source in terms of discrete lexical material or case, person, number, and/or TAM markings while otherwise maintaining the lexical and structural content needed to appropriately translate the source text.
@@ 5.1 Quality metrics
We present BLEURT scores for prompts generated using automated retrieval in Table 1 and summarize human quality judgments for GPT-3.5 and GPT-4o with automated retrieval in Table 2. The complete distribution of quality ratings across all prompt types for these two models is provided in Appendix F. We find clear effects of LLM, prompt type, and retrieval method, as well as interactions among all three factors.
Gemini and GPT-4o outperform Llama 3 and GPT-3.5 for every prompt type. This gap is highest for the least informative prompts, indicating that the Llama 3 and GPT-3.5 base models have relatively poor coverage of S. Quechua, while GPT-4o and Gemini have much better coverage. The effect is evident in both BLEURT scores and human quality evaluations.
@@ 5.4 Mistranslations
We also note that many of the items tagged as Mistranslation with Lexical Correspondence show correspondence only for words that were already in Spanish in the source text. For example, some sentences contain Spanish loan names for the days of the week. While some of these errors are produced using deceptively fluent Spanish as described above, we find many to be accompanied by semantic incoherence or ungrammaticality in the output. We discuss such target language fluency errors in the following section.
@@ 5.6 Grammatical Divergence
We group misrendered verbal subjects and tense/aspect/morphology (TAM) markers together as Grammatical Divergence errors. Such errors are distinct from the Target Fluency errors described in the previous section— the Spanish output is grammatical, but fails to accurately reflect the content from the source. Although they are not the only grammatical phenomena that may be similarly misrendered, we select subject and TAM markers for analysis as they are straightforward to identify and give a good indication of how well the LLMs cope with more abstract information about the meanings of functional morphemes. TAM divergences are much more prevalent than divergences in subject; for instance, only one of GPT-4o’s 13 Grammatical Divergence errors in the morph condition misrender the subject marker.

### v2

@@ Abstract
This work investigates the in-context learning abilities of pretrained large language models (LLMs) when instructed to translate text from a low-resource language into a high-resource language as part of an automated machine translation pipeline. We conduct a set of experiments translating Southern Quechua to Spanish and examine the informativity of various types of context retrieved from a constrained database of digitized pedagogical materials (dictionaries and grammar lessons) and parallel corpora. Using both automatic and human evaluation of model output, we conduct ablation studies that manipulate (1) context type (morpheme translations, grammar descriptions, and corpus examples), (2) retrieval methods (automated vs. manual), and (3) model type. Our results suggest that even relatively small LLMs are capable of utilizing prompt context for zero-shot low-resource translation when provided a minimally sufficient amount of relevant linguistic information. However, the variable effects of context type, retrieval method, model type, and language-specific factors highlight the limitations of using even the best LLMs as translation systems for the majority of the world’s 7,000+ languages and their speakers.
@@ 1 Introduction
Despite great progress in the quality of today’s state of the art machine translation (MT) systems, constraints on the amount and kinds of data available in the majority of the world’s 7,000+ languages have led to yet another disparity in access and support for speakers of these languages: low-resource MT continues to be a major challenge Hendy et al. (2023); Nicholas and Bhatia (2023); Robinson et al. (2023); Stap and Araabi (2023). Although many languages lack the kinds of large, standardized corpora necessary for traditional MT methods, recent work suggests it may be possible to leverage a smaller amount of existing resources, for example pedagogical materials used for language instruction, to develop MT systems with Large Language Models (LLMs), albeit with varying results Elsner and Needle (2023); Tanzer et al. (2024); Zhang et al. (2024). These materials are often the result of community-driven or government-led initiatives to support language revitalization, reclamation, and mother-tongue education Schreiner et al. (2020); Liu et al. (2022); Riestenberg et al. (2024). Such discrepancies in the needs and priorities of academic, commercial, and community-led efforts to develop digital resources and language technologies is what Gessler (2022) terms the “NLP Gap”.
@@
Our results suggest that while, unsurprisingly, translation quality improves with model size, such improvements seem to primarily be the result of previous exposure to the low-resource language during model pretraining, rather than an improved ability for the model to utilize prompt context, as evidenced by high scores in response to baseline (zero-shot) translation prompts. However, we also find evidence that in-context learning abilities may be inconsistent across different models of similar size. As found in previous studies, prompts containing morpheme and word-level translations reliably improve model outputs, but information from the grammar and corpus have a null or even negative effect on results. Human evaluation on a selection of outputs from two models – GPT-3.5 Turbo and GPT-4o – align with the quantitative measures we obtain using BLEURT Sellam et al. (2020) as an automatic metric. Quantitative results also show an effect of automated retrieval on translation quality that is most evident in prompts containing morpheme translations and for models with lower baseline scores. Finally, we highlight a number of ethical concerns and limitations that arise from the proposed methods that are supported by our findings, and discuss the potential risks and challenges LLM-based methods for low-resource MT face moving forward.
@@ 2 LLMs for Machine Translation
Zhang et al. (2024) find that sentences from the corpus retrieved using BM25 embeddings Robertson et al. (2009) work better than random ones. Tanzer et al. (2024), however, report that retrieval with longest common substring (LCS) matching outperforms embedding-based retrieval. Overall, the question of how to best retrieve relevant passages containing grammar material or sentences in a low-resource language is still open. This also complicates the interpretation of the mostly-negative results found for grammar passages. It is not clear whether these stem from poor retrieval, from the LLM’s inability to process the retrieved content, or both. Moreover, although Reid et al. (2024) conducts human evaluation of the results for quality, to the best of our knowledge no study to date systematically investigates specific grammatical errors in the output.
Finally, each of these studies finds a significant decrease in LLMs’ abilities to translate from a high-resource language into a low-resource language relative to experiments in the opposite direction. This is in line with McCoy et al. (2023), who find that while the accuracy of an LLM’s output highly depends on the probability of both the input and the output text, output probability has a greater influence on model performance. We therefore focus this study on a single translation direction, instructing LLMs to output translations from a low-resource language into the language with which they are likely to have had more exposure during training, i.e, from Southern Quechua into Spanish, and leave the reverse direction for future work.
@@ 3 Quechuan Languages
Quechua is a family of languages Indigenous to the Andes in South America. This study focuses on varieties of Southern Quechua (S. Quechua, also known as urin quechua or quechua sureño) spoken in parts of Peru. [^1] Unless noted otherwise, we use Quechua in this study to refer Southern Quechua and related varieties. While previous studies investigated language-LLM pairs for which the baseline LLM lacked any pretrained knowledge, we find that newer LLMs can translate some S. Quechua sentences in a zero-shot setting. We expect this to be typical of many low-resource languages which, while often endangered, still may have some presence on the web.
@@
Since Quechua is primarily spoken in South America, the majority of available digital resources, including all materials used in this study, use Spanish as the language of translation, explanation, and/or instruction. We therefore also use Spanish, rather than English or any other high-resource language, as the language of translation and prompting when testing our system.
3.1 Language-Specific Factors
While the proposed methods are general enough to be applied to any language pair, model outputs may reflect certain language-specific characteristics of the source and target languages, respectively. In this section, we provide a brief description of selected language-specific factors in S. Quechua as they relate to their translated Spanish counterparts. For a discussion of their potential effects on our results, please see Section 6.1.
3.1.1 Morphological Segmentation
S. Quechua is primarily agglutinating, i.e., much of the morphology may be described in terms of isomorphic form-meaning relationships, morphemes generally maintain a consistent form regardless of their phonological environment, and morpheme boundaries tend to be transparent. In contrast, morpheme segmentation in Spanish may be rendered opaque due to its fusional morphology and widespread use of conditional allomorphy.
While LLMs are trained to process text via token-based rather than morpheme-based segmentation, it is possible that a lack of direct correspondence between the expression of morphosyntactic categories in S. Quechua and Spanish may affect a model’s ability to leverage the information we provide as prompt context in our experiments. Correspondences in form and meaning across parallel usage examples may be particularly obscured, limiting the use of corpora designed for traditional MT methods. It may be possible to mitigate such issues with more advanced retrieval or prompting techniques, for example by explicitly instructing an LLM to conduct morphological analysis as part of the translation process, but we leave this for future work.
3.1.2 Syncretism and Polysemy
Although the language is primarily agglutinating, a number of morphemes in S. Quechua are syncretic, such that a given form may be used to express more than one grammatical category. For example, the 1sg.poss marker, -y, shares the same form as both the 2sg.imp marker and the infinitival marker, as illustrated in the following examples:
[…]
. ñaña-y
sister-1sg.poss
‘my sister’
[…]
. Mikhu-y!
eat-2sg.imp
‘Eat!’
[…]
. mikhu-y-ta muna-ni
eat-inf-acc want-1sg
‘I want to eat’
Similarly, words in S. Quechua may be polysemous, with the potential to express more than one meaning depending on their use in context. For example, the S. Quechua word miski (misk’i) may be translated as either dulce ‘sweet’ or rico/a ‘delicious’, and both dulce and rico/a are themselves polysemous in Spanish. Dulce may be used as an adjective, i.e., ‘sweet’, or a noun, i.e.,‘candy’, and rico/a may describe either richness in flavor, i.e., ‘delicious’, or in monetary wealth, i.e., ‘rich’.
The exact forms displaying syncretism or polysemy must be identified on a language-specific basis, but the ambiguity they present poses a clear problem for our proposed methods in general, with potential effects on both retrieval and generation. We discuss this issue further in Section 6.1.
3.1.3 Variation
Both S. Quechua and Spanish are characterized by extensive regional and dialectal variation. In S. Quechua, this includes differences in orthographic and/or phonological conventions as well as the specific lexical items and expression of morphosyntactic content. For example, the S. Quechua word for ‘dog’ may be rendered orthographically as alqo, allqo, allku, allqu, or ashko, and the additive suffix may be expressed as either -pas or -pis, depending on the community. Variation in the attested usage of specific lexical items and morphemes across communities is also common in S. Quechua. For example, the evidential marker -mi /-m is frequently attested in the Peruvian variety of S. Quechua used in this study, but essentially absent in many Bolivian varieties.
Variation across Spanish-speaking communities may also affect models’ abilities to produce translations that are both accurate and appropriate. The Andean Spanish reference translations used in this study do not appear to affect the results of our automatic evaluation. However, were the proposed methods to be applied in a realistic setting, it would be especially important to assess the degree of alignment between any prescriptive linguistic standards that have been implicitly acquired by the LLM and the usage conventions of the language community or communities of interest.
@@ 4.1 Data
We conduct experiments on a collection of 50 pairs of S. Quechua - Spanish sentences sourced from one of the author’s personal notes. These were selected to highlight a range of specific grammatical phenomena at multiple levels of difficulty— they include simple clauses and tenses (Example 4.1), as well as more advanced constructions such as those involving past participles (Example 4.1) and simultaneous events (Example 4.1.
[…]
. qam allin-ta tusu-nki
you good-acc dance-2sg
tu bailas bien
‘you dance well’
[…]
. awa-sqa-y wali-qa sumaq-mi
knit-pp-1sg skirt-top great-assertive
la falda que tejí es linda
‘the skirt that I knit is pretty’
[…]
. qam-qa taki-ta uyari-spa wasi-yki-ta picha-chka-nki
you-top song-acc listen-subr house-2.poss-acc clean-prog-2sg
tú estás limpiando tu casa escuchando música
‘you’re cleaning your house listening to music’
The first author, a foreign-language student of S. Quechua, received permission from her instructor to use notes from their lessons for the study. All sentence pairs were inspected by the instructor, a native bilingual speaker of both S. Quechua and Peruvian Spanish, to eliminate any errors and confirm the accuracy of all reference translations.
@@ 4.2 Prompt Construction
[TAREA] Traduce la siguiente frase del quechua al español. Responde sólo con la traducción: quechua: kay wasiqa turiypam español:
Figure 1: Example baseline prompt. English: [TASK] Translate the following sentence from Quechua to Spanish. Respond only with the translation: Quechua: kay wasiqa turiypam; Spanish:
@@ 4.2.1 Morpheme Translations (morph)
We use a morphological parser Rios (2015) to segment each word of the source segment into morphemes, each with gloss symbols and a Spanish translation. [^2] We set aside valid concerns regarding the theoretical status of the morpheme for this study and define a morph(eme) loosely as a recognizable form-meaning pair that recurs in a language. Some morphemes have multiple candidate meanings, all of which are retrieved. As an example, the word rantikuq is segmented as ranti-ku-q and glossed as “comprar.DB.VRoot-DB.VDeriv.+RflxInt+Ag.NS.”
While numerous orthographic standards have been developed and promoted across Quechuan-speaking communities in South America, considerable variation in orthographic conventions may be found even within a particular community or variety Rios and Castro Mamani (2014). We discuss the implications of this for our results in Section 4.2.5.
We supplement the output from the parser using a Quechua-Spanish bilingual dictionary Qheswa Simi Hamut’ana Kurak Suntur (2005). We retrieve any dictionary entry whose headword exactly matches a morpheme in our segmentation. By default, we include all senses and any usage examples or contextual information in the dictionary entry as part of the prompt. We then concatenate the output of the parser with the retrieved dictionary entries and include this morph information as prompt context preceding the source sentence and baseline translation prompt.
@@ 4.2.2 Grammar Descriptions (grammar)
We also experiment with the inclusion of grammar lessons found in student-facing pedagogical materials, retrieving grammatical explanations relevant to each source sentence from a PDF document developed for students and teachers of S. Quechua. Pinto Tapia et al. (2005). The document is organized into short sections (1-3 sentences, plus paradigm tables or usage examples) that describe the particular grammatical concept associated with an affix in Quechua. For each source sentence, we retrieve sections associated with any affix listed in the document that is an exact match of a morpheme and include this in prompts using contextual information from the grammar. This improves on the methods described in Tanzer et al. (2024), who use LCS-based retrieval over an entire textbook, and Elsner and Needle (2023); Zhang et al. (2024), whose grammatical description remains consistent across prompts regardless of the source text being translated.
@@ 4.4 Evaluation
We conduct both automatic and human evaluations to identify trends in model errors and outputs in the various experimental conditions. We calculate BLEURT and BLEU scores as automatic metrics, and report mean BLEURT scores across items as the primary quantitative measure of translation quality for each of the conditions and models. We also use an adapted MQM schema to conduct qualitative human evaluation of the outputs of GPT-3.5 and GPT-4o for all prompts using automatic retrieval.
Each item selected for human evaluation is annotated by at least one of the authors by comparing the model’s output to the source text and reference translation. We refer to the complete MQM typology to design our own four-dimensional framework of commonly attested errors in LLM-MT, each with a defined set of specific subtypes. Precise definitions and examples for all error categories and subtypes may be found in Appendix D.
Many of the categories in our schema are defined as in the core MQM framework. However, to capture some of the key behaviors reported in previous studies on LLM-MT and to evaluate the effects of prompt type on model outputs, we make the following adjustments. First, we utilize the Addition and Omission errors defined as Accuracy subtypes in the original MQM typology, but distinguish these from three additional subtypes: Substitution - Incorrect Subject, Substitution - Incorrect Tense/Aspect/Modality (TAM), and Substitution - Other. This is intended to capture LLM translations that differ from the source in terms of discrete lexical material or case, person, number, and/or TAM markings while otherwise maintaining the lexical and structural content needed to appropriately translate the source text. Although they are not the only grammatical phenomena that may be similarly misrendered, we select subject and TAM markers for analysis as they are straightforward to identify and give a good indication of how well the LLMs cope with more abstract information about the meanings of functional morphemes.
@@ 5.1 Quality Metrics
We present BLEURT scores for prompts generated using automated retrieval in Table 1 and summarize human quality judgments for GPT-3.5 and GPT-4o with automated retrieval in Table 2. The complete distribution of BLEURT, BLEU, and human-annotated quality ratings for all of our experiments is provided in Appendix F. We find clear effects of LLM, prompt type, and retrieval method, as well as interactions among all three factors.
Comparing across models, we find that Gemini and GPT-4o outperform Llama 3 and GPT-3.5 for every prompt type. This gap is highest for the least informative prompts, indicating that the Llama 3 and GPT-3.5 base models have relatively poor coverage of S. Quechua, while GPT-4o and Gemini have much better coverage. The effect is evident in both automatic and human quality evaluations.
@@ 5.4 Mistranslations
We also note that many of the items tagged as Mistranslation with Lexical Correspondence show correspondence only for words that were already in Spanish in the source text. For example, some sentences contain Spanish loan names for the days of the week. While some of these errors are produced in deceptively fluent Spanish, we find many to be accompanied by semantic incoherence or ungrammaticality in the output. We discuss such target language fluency errors in the following section.
@@ 5.6 Grammatical Divergence
We group misrendered verbal subjects and tense/aspect/morphology (TAM) markers together as Grammatical Divergence errors. Such errors are distinct from the Target Fluency errors described in the previous section— the Spanish output is grammatical, but fails to accurately reflect the content from the source. TAM divergences are much more prevalent than divergences in subject; for instance, only one of GPT-4o’s 13 Grammatical Divergence errors in the morph condition misrender the subject marker.